OpenStudy (luigi0210):

Find dy/dx of:

OpenStudy (luigi0210):

Y= integral of 1/t of dt where x>o from 1 to x

OpenStudy (anonymous):

you mean: \[\frac{d}{dx}\left[ \int\limits \frac{dt}{t}\right]\]

OpenStudy (anonymous):

sry, forgot integration limits: \[\frac{d}{dx}\left[\int\limits_0^x\frac{dt}{t}\right]\]

OpenStudy (anonymous):

this?

OpenStudy (anonymous):

if that's it, then dy/dx=1/t, by mane theorem of calculus

OpenStudy (luigi0210):

No, I mean dy/dx of y= (integral) (upper limit is 1, lower limit is x) then the equation of (1/t) with respect to t Sorry I can't do an integral sign from where I'm typing this

OpenStudy (luigi0210):

And oh

OpenStudy (anonymous):

\[\int\limits_x^1=-\int\limits_1^x\] so answer is -1/t